### v1

# Eigendecompositions of temporal networks

Abstract
Temporal networks, defined as sequences of time-aggregated adjacency matrices, sample latent graph dynamics and trace trajectories in graph space. By interpreting each adjacency matrix as a different time snapshot of a scalar field, fluid-mechanics theories can be applied to construct two distinct eigendecompositions of temporal networks. The first builds on the proper orthogonal decomposition (POD) of flowfields and decomposes the evolution of a network in terms of a basis of orthogonal network eigenmodes which are ordered in terms of their relative importance, hence enabling compression of temporal networks as well as their reconstruction from low-dimensional embeddings. The second proposes a numerical approximation of the Koopman operator, a linear operator acting on a suitable observable of the graph space which provides the best linear approximation of the latent graph dynamics. Its eigendecomposition provides a data-driven spectral description of the temporal network dynamics, in terms of dynamic modes which grow, decay or oscillate over time. Both eigendecompositions are illustrated and validated in a suite of synthetic generative models of temporal networks with varying complexity.
Network science [1] models how individual elements interact with each other in a complex system, and how such architecture backbone conditions and influences the emergence of complex dynamical phenomena. When interactions also fluctuate over time, networks (or graphs) are themselves time-varying, leading to the theory of temporal networks [2, 3]. Among other possible definitions, in this work we call a Temporal Network (TN) a (time) ordered sequence of graphs. While complexity science has predominantly investigated how dynamics defined on top of network structures are affected when such structures are time-varying [4, 5, 6, 7, 8, 9], lesser work has considered principled analysis of the intrinsic dynamics of temporal networks [10]. In this regard, a recent approach considers TNs as time series of graphs –i.e. trajectories in graph space of a possibly latent graph dynamics–, paving the way for extensions to the network realm of classical dynamical systems and time series analysis concepts such as temporal autocorrelation [11, 12, 13, 14] or Lyapunov exponents [15, 16, 17].
Here we take this latter view and further interpret the sequence of adjacency matrices which defines a TN as sampling a discrete scalar field whose evolution is given by a latent graph dynamics [18]. This, we shall see, enables to use some tools from fluid mechanics [19] to characterise such network trajectories and to approximate the latent graph dynamics. Building on two different fluid-mechanic methodologies –namely proper orthogonal decomposition (POD) [20] and dynamic mode decomposition (DMD) [21], the latter being one possible numerical implementation of the Koopman operator [22, 23, 24]– we propose two mathematically principled eigendecompositions of temporal networks, whose properties can be leveraged, respectively, for compressing the TN’s structure and finding low-rank embeddings [25, 26], and for characterizing, forecasting and controlling its latent dynamics [27, 28].
Theory – A temporal network (TN) is defined as an ordered sequence of $m$ real-valued adjacency matrices $({\bf A}(1),{\bf A}(2),\dots,{\bf A}(m))$, where ${\bf A}({t})\in\mathbb{R}^{n\times n}$ aggegates the activity of a latent graph dynamics between t1t-1 and $t$. Let us suitably flatten each adjacency matrix into a vector
${\bf a}(t)=\texttt{flat}({\bf A}({t})-\langle{\bf A}\rangle),$ (1)
where $\langle{\bf A}\rangle=\frac{1}{m}\sum_{t=1}^{m}{\bf A}(t)$ is the time-averaged adjacency matrix (sometimes called the annealed adjacency matrix), and flattening is such that $\texttt{flat}({\bf A})=(A_{11},A_{12},\dots,A_{1n},A_{21},A_{22},\dots,A_{2n},\dots,A_{n1},\dots,A_{nn})$. Accordingly, ${\bf a}(t)\in\mathbb{R}^{n^{2}}$ is a real-valued column vector [^1] For simple, undirected networks, $n^{2}$ reduces to n(n1)/2n(n-1)/2, and for unweighted graphs, entries are binary.. By subsequently stacking each of the $m$ flattened vectors in order, the TN admits a $n^{2}\times m$ matrix representation $\mathscr{S}$} such that
${\mathscr{S}}\equiv{\mathscr{S}}_{[1\dots m]}:=[{\bf a}(1)\ {\bf a}(2)\ \dots\ {\bf a}(m)].$ (2)
In this way, each network snapshot is nothing but a point in a $n^{2}$-dimensional Euclidean space.
Our first goal is to be able to project each (normalized) network snapshot into a space spanned by a base of only $r$ (with $r\ll n^{2}$) orthogonal network modes $\{\phi_{j}\}_{j=1}^{r}$ such that
${\bf a}(t)\approx\sum_{j=1}^{r}\alpha_{j}(t)\phi_{j},\ t=1,2,\dots,m,$ (3)
where $\phi_{j}\in\mathbb{R}^{n^{2}}$ are some specific directions in graph space characterizing a suitable new base and $\boldsymbol{\alpha}(t)=[\alpha_{1}(t),\alpha_{2}(t),\dots,\alpha_{r}(t)]$ are the coordinates of the $t$-th snapshot in this lower-dimensional space, respectively. Observe that Eq. 3 trivially becomes an identity when $\phi_{j}$ are the canonical unit vectors in $\mathbb{R}^{n^{2}}$ and $r=n^{2}$, or when $\phi_{j}$ coincide with the eigenvectors of the covariance matrix ${\mathscr{S}}{\mathscr{S}}^{\top}\in\mathbb{R}^{n^{2}\times n^{2}}$. In this latter case, each $\phi_{j}$ is a suitable linear combinations of the canonical unit vectors. After de-flattening, $\{\phi_{j}\}_{j=1}^{n^{2}}$ form a set of concrete adjacency matrices that we call network eigenmodes. In the realm of fluid mechanics, this latter decomposition –applied to suitably flattened flowfield snapshots– is called proper orthogonal decomposition (POD) [20, 19], and decomposes the fluid solution of some flow equation –e.g. Navier-Stokes-equations over a fixed domain– in terms of a suitable combination of spatial coherent structures. Incidentally, POD shares similarities with other equally popular concepts across the disciplines like the Karhunen–Loeve decomposition in signal processing or principal component analysis (PCA) in data science [29], and traces back to the theory of singular value decomposition [30]. In the context of network theory, very recently Thongprayoon et. al proposed an analogous approach to use PCA for tie-decay temporal network embedding [25]. As a matter of fact, the eigenvectors $\phi_{j}$ also provide the optimal [^2] In the $L_{2}$ norm. low-rank truncation of $\mathscr{S}$} [30, 31], i.e. for a given $r<n^{2}$, one cannot approximate better ${\bf a}(t)$ than with Eq. 3 where $\phi_{j}$ are the network eigenmodes. Furthermore, since covariance matrices are symmetric and semi-positive definite, these network eigenmodes $\{\phi_{j}\}$ can be ordered in terms of their associated eigenvalue $\{\lambda_{j}\}$ –which indicates the relative energy present in the mode–, and the spectral decomposition of $\mathscr{S}\mathscr{S}^{\top}$ that reads $\mathscr{S}\mathscr{S}^{\top}=\sum_{i=1}^{n^{2}}\lambda_{i}\phi_{i}\phi_{i}^{\top}$ can be ordered such that $\lambda_{1}\geq\lambda_{2}\geq\cdots\geq\lambda_{n^{2}}\geq 0$. Accordingly, the $r$ orthogonal directions in graph space coincide with the network eigenmodes with larger associated eigenvalue. Finally, the coordinates of each network snapshot in the new, $r$-dimensional space are found by projecting each network snapshot into the basis of modes
$\alpha_{j}(t)=\langle{\bf a}(t_{j}),\phi_{j}\rangle.$ (4)
Incidentally, observe that when the number of nodes $n$ in each network snapshot is typically larger than the temporal resolution (more concretely, when $n^{2}\gg m$), computationally speaking it is wiser to compute the eigendecomposition of the covariance matrix of $\mathscr{S}$$^{\top}$, i.e. ${\mathscr{S}}^{\top}{\mathscr{S}}=\sum_{i=1}^{m}\lambda_{i}\psi_{i}\psi_{i}^{\top}$, because the covariance matrices ${\mathscr{S}}{\mathscr{S}}^{\top}\in\mathbb{R}^{n^{2}\times n^{2}}$ and ${\mathscr{S}}^{\top}{\mathscr{S}}\in\mathbb{R}^{m\times m}$ have the same set of non-null eigenvalues $\{\lambda_{i}\}_{i=1}^{m}$ ($\mathscr{S}$} has at most rank $m$), and their associated eigenvectors $\{\phi_{i}\}_{i=1}^{n^{2}}$ and $\{\psi_{i}\}_{i=1}^{m}$ are related via
$\phi_{j}=\frac{1}{\sqrt{\lambda_{j}}}{\mathscr{S}}\psi_{j},\forall j\ \text{s.t.}\ \lambda_{j}\neq 0.$
In a nutshell, the $m$ snapshots of a temporal network –which originally live in a space as large as the number of entries of the adjacency matrix– can be optimally projected onto an $r$-dimensional space by leveraging on the eigendecomposition of a suitable temporal network’s covariance matrix. The reconstruction error of the TN by using the compressed representation is controlled by the cumulated sum of the first $r$ eigenvalues.
Our second goal is to apply ideas from Koopman operator theory to the realm of temporal networks, much as dynamic mode decomposition (DMD) [21] applies Koopman theory to fluid mechanical problems [19]. Given a –possibly nonlinear– finite-dimensional dynamical system acting on a finite state space $\mathscr{X}$ (e.g. the latent graph dynamics which, adequately sampled, generates a temporal network formed by a sequence of weighted adjacency matrices), the Koopman operator is a linear operator acting on a (possibly infinite-dimensional) observable $g(\mathscr{X})$, such that $g(\mathscr{X}[t+1])=\mathscr{K}g(\mathscr{X}[t])$. Koopman theory indeed proves that such operator –in general, defined in a Hilbert space– always exist, but it does not prescribe a recipe to find the adequate observable [^3] Observe that when the observable is identified with the state density –i.e. the marginal distribution of states–, the adjoint of the Koopman operator coincides with the transfer (Perron-Frobenius) operator., and in complex systems it is in general challenging to discover the coordinate transformations (observable) that linearize the dynamics [32].
Following the dynamic mode decomposition technique [21, 19], the simplest way to numerically approximate the (possibly infinite-dimensional, linear) Koopman operator $\mathscr{K}$ to the latent graph dynamics generating the temporal networks is probably to assume a linear evolution in the actual, finite state space ${\bf a}\in\mathbb{R}^{n^{2}}$, i.e. $g(\mathscr{X})={\bf a}$. In this case, the operator is the best-fit matrix $\mathscr{K}\in\mathbb{R}^{n^{2}\times n^{2}}$ that advances the network trajectory, i.e. such that ${\bf a}(t+1)\approx\mathscr{K}{\bf a}(t),\forall t=1,2,\dots,m$. From Eq. 5, we define two overlapping $\mathbb{R}^{n^{2}\times m-1}$ matrices
${\mathscr{S}_{1}}={\mathscr{S}}_{[1\dots m-1]}=[{\bf a}(1)\ {\bf a}(2)\ \dots\ {\bf a}(m-1)];$
${\mathscr{S}_{2}}={\mathscr{S}}_{[2\dots m]}=[{\bf a}(2)\ {\bf a}(3)\ \dots\ {\bf a}(m)],$ (5)
such that least-squares fitting $\mathscr{K}$ is the (low-rank [^4] Imposing low-rank is just one possible restriction. In the realm of quantum mechanics, we instead impose $\mathscr{K}$ to be a self-adjoint operator. In general, other physics-informed schemes can impose different structures. For instance, if energy is to be conserved, then $\mathscr{K}$ must be a unitary operator (reflection or rotation). In our case, if e.g. if networks are unweighted and the number of links is preserved along the dynamics, then $\mathscr{K}$ is a permutation matrix, i.e. we impose it to be an orthogonal binary matrix.) solution of
[…]
Since $\mathscr{K}$ is a linear operator, its eigenvectors correspond to coherent network structures (often called dynamic modes) and the eigenvalues describe their stylized growth, decay or oscillation that, all combined, characterise the linear approximation to the latent graph dynamics. Observe that if such latent dynamics is highly nonlinear, choosing $g(\mathscr{X})={\bf a}$ may yield a poor approximation to the true Koopman operator (e.g. time forecasting becomes a challenge). Accordingly, in that case one needs to look for a more sophisticated $g(\cdot)$ on which the linear action of $\mathscr{K}$ takes place. Invoking Takens embedding theorem [^5] This theorem states that the attractor of high-dimensional nonlinear dynamics is diffeomorphic to the one reconstructed from any one-dimensional time series projection of the system, provided such projection is adequately embedded in sufficiently large time-delayed coordinates. [33, 34] one possible choice for such observable is to build a time-delay embedding of ${\bf a}(t)$, i.e. $g(\mathscr{X})=\tilde{{\bf a}}(t)=[{\bf a}(t),{\bf a}(t-1),\dots,{\bf a}(t-d)]$, where $d$ is the number of additional, time-delayed coordinates. Under this choice, the matrix-form of the trajectories $\tilde{\mathscr{S}}_{X}=[\tilde{{\bf a}}(1)\ \tilde{{\bf a}}(2)\ \dots\ \tilde{{\bf a}}(m-1)]$ is Hankel [35, 36]. While the $d=0$ case is intuitively simpler, note that the more general case $d>0$ is not really much more computationally harder, since while the dimensionality of $\mathscr{K}$ increases from $\mathbb{R}^{n^{2}\times n^{2}}$ to $\mathbb{R}^{n^{2}(1+d)\times n^{2}(1+d)}$, $\mathscr{K}$ is actually never explicitly computed nor diagonalized as we see below.
By construction, we have $\mathscr{K}\approx\mathscr{S}_{2}\mathscr{S}_{1}^{+}$, where $\mathscr{S}_{1}^{+}$ is the Moore-Penrose pseudoinverse of $\mathscr{S}_{1}$. The linear dynamics of the TN is thus fully described by the eigendecomposition of $\mathscr{K}$, hence providing a data-driven way of assessing the intrinsic dynamics of temporal networks using the tools of linear dynamics. Now, instead of directly computing and then diagonalizing $\mathscr{K}$ –usually an ill-conditioned, and computationally hard problem for large networks, both for simple observables or time-delayed coordinate embedddings–, we leverage singular value decomposition (SVD) theory to compute the leading terms of its spectral decomposition as follows. First, ${\mathscr{S}}_{1}$ admits an SVD ${\mathscr{S}}_{1}=U\Sigma V^{*},$ where, incidentally, the columns of $U$ are nothing but the network eigenmodes previously discussed in Eq. 3. Similarly, ${\mathscr{S}}_{2}=\mathscr{K}U\Sigma V^{*}$. By further projecting into the network eigenmodes, we then have
$U^{*}{\mathscr{S}}_{2}V\Sigma^{-1}=U^{*}\mathscr{K}U:=\tilde{\mathscr{K}}$, (7)
where $\tilde{\mathscr{K}}\in\mathbb{R}^{m\times m}$ is in general much smaller than $\mathscr{K}$. $\tilde{\mathscr{K}}$ describes how the network eigenmodes evolve in time through the action of the linear dynamics. We then diagonalize $\tilde{\mathscr{K}}=\sum_{i}\Lambda_{i}{\bf w}_{i}{\bf w}_{i}^{\top}$, where $W=[{\bf w}_{1}\ {\bf w}_{2}\ \dots]$ is eigenvector matrix and $\Lambda_{i}$ is the eigenvalue associated to ${\bf w}_{i}$. Now, by virtue of the properties of the SVD, the non-zero eigenvalues of $\mathscr{K}$ coincide with the eigenvalues of $\tilde{\mathscr{K}}$. The matrix containing the eigenvectors associated to the non-null eigenvalues of $\mathscr{K}$, which we denote $\Phi$, is finally calculated as [21]
$\Phi=\mathscr{S}_{2}V\Sigma^{-1}W$
(incidentally, when $\mathscr{S}$ admits a low-rank approximation (i.e. $r\ll\min(n^{2},m)$), we use the reduced SVD instead of the full SVD). Finally, the growth/decay of each eigenvector or their frequency of oscillation corresponds, respectively, to the real and imaginary parts of $\log(\Lambda_{i})$: eigenvalues inside the unit circle denote stable (exponentially decaying) modes, those outside the unit circle denote unstable (exponentially growing) modes, and eigenvalues which are roots of one yield pure oscillation with no growth or decay.
Figure 1: (A,B,D,E) The first two network eigenmodes obtained from the eigendecomposition of the covariance matrix $\mathscr{S}\mathscr{S}^{\top}$, for asynchronous, for a temporal network of $m=100$ network snapshots of $n=50$ nodes, with asynchronous, noisy sinusoidal link dynamics of the form $A_{ij}(t)=\cos(2\pi t/T_{1}+\eta_{ij})+\cos(2\pi t/T_{2}+\eta_{ij})+\xi$, with two concomitant periods $T_{1}$ and $T_{2}$ (see the text for details). The first two eigenmodes concentrate about $80\%$ of the variance. (C,F) Projection of the temporal network trajectory onto the space spanned by the first two network eigenmodes, resulting in a periodic orbit with period LCM($T_{1},T_{2}$).
Figure 2: Temporal network projection into the two-dimensional space spanned by the first two network eigenmodes, for TNs of $n=30$ nodes and $m=500$ time steps, for (A) a white temporal network, (B) a Random Walk TN, (C) a DARN(1) process, (D) a DARN(12) process, and (F) a low-dimensional chaotic network (dictionary trick). The energy stored in the first two eigenmodes amounts to (A) $2.5\%$, (B) $50\%$, (C) $10\%$, (D) $19\%$ and (F) $80\%$. (E) Energy (variance) stored in the first two eigenmodes for DARN(p) TNs, as a function of the memory order pp.
Results – We have constructed generative temporal networks with a suite of different intrinsic dynamics, including (i) white noise TNs, (ii) different types of noisy periodic dynamics, (iii) autorregressive network processes with memory, and (iv) network chaotic dynamics, see [26, 11, 15] for details. Results overall indicate that the intrinsic dynamics of the TN is systematically well preserved in the orbits projected in the two-dimensional plane spanned by the first two network eigenmodes, yielding a substantial TN compression for all models except for white noise TN, which by construction cannot be compressed. The coordinates of these two-dimensional orbits are $(\alpha_{1}(t),\alpha_{2}(t))$ and are computed from Eq. 4. At the same time, note that the actual TN dynamics can be reconstructed from the compressed system simply using Eq. 3.
As an illustration, consider a TN built by concatenating several times a block of $T$ Erdos-Renyi network snapshots, so that the TN has period $T$ and is fully uncorrelated at lags different than $T$ or its multiples [11]. After computing the eigendecomposition of $\mathscr{S}\mathscr{S}^{\top}$, we find that the first $T-1\ll n^{2}$ network eigenmodes concentrate about $100\%$ of the variance. Furthermore, the trajectory of the TN projected in the first two network eigenmodes is itself a periodic orbit of $T$ points, yielding a massive compression with small reconstruction error.
Subsequently, we build a different periodic TN model, where the adjacency matrices have asynchronous and multiple-period link dynamics $A_{ij}(t)=\cos(2\pi t/T_{1}+\eta_{ij})+\cos(2\pi t/T_{2}+\eta_{ij})+\xi$, where $T_{1}$ and $T_{2}$ are the two intrinsic periods, $\eta_{ij}\sim\textsc{uniform}(0,1)$ is a quenched noise that desynchronizes the dynamics of each link, and $\xi\sim\textsc{N}(0,\sigma)$ is a Gaussian dynamical noise. Fig. 1 displays the first two network eigenmodes (panels A and B, D and E) and the projection into these eigenmodes (panels C and F) of a TN of $m=100$ time snapshots and $n=50$ nodes following this dynamics with $\sigma=0.5$ and $T_{1}=20,T_{2}=10$ (panel C) and $T_{1}=6,T_{2}=4$ (panel F). The periodic nature of the TN is again well captured by the orbit in the 2-dimensional network eigenmode space, which is itself quasi-periodic (periodic for noise $\sigma\to 0$) with a quasi-period equal to the lowest common multiple (LCM) of $T_{1}$ and $T_{2}$ ($\text{LCM}(20,10)=20,\text{LCM}(6,4)=12$).
As a complement, in Fig. 2 we plot the two-dimensional orbits in eigenmode space associated to temporal networks with $n=20$ nodes, $m=500$ snapshots and generated by (A) a white Erdos-Renyi temporal network (a sequence of i.i.d. ER graphs, lacking structure), (B) a random walk network trajectory, (C-D) a discrete autorregressive network DARN($p$) [10] with memory order $p=1$ (C) and $p=12$ (D) respectively, where each link $\ell_{t}$ is constructed independently, such that with probability $q=0.9$, $\ell_{t+1}$ samples uniformly from its past $p$ states, and with probability (1q)(1-q), it assigns a Bernoulli trial with probability $y=0.5$ [11, 15, 26]. Trajectories in eigenmode space mimic random walks with increasing memory for increasing $p$. (E) Total variance explained by the first two network eigenmodes, as a function of the memory order $p$ of the DARN($p$) process, showing that the embedded trajectory is less random for increasing memory orders, as expected. (F) A low-dimensional chaotic network generated by the dictionary trick on the fully chaotic logistic map [11, 15], where the low-dimensional map becomes evident in the eigenmode space.
Figure 3: Description of Conway’s Game of Life 2D Cellular Automaton, interpreted as a temporal network of $n=50$ nodes. (A) Evolution of the 2DCA in the space spanned by the first two network eigenmodes (depicted in panel (C)). This trajectory is similar to a random walk with memory, see Fig. 2C. (B) Mean-square displacement, showing superdiffusive behavior.
As an application, we identify the subsequent states of 2D cellular automata (2DCA) as the adjacency matrices of a temporal network. We implement Conway’s Game of Life [37] –a Turing-complete 2DCA [38]– for a grid of $20\times 20$ (i.e., $n=20$ nodes) and $300$ evolution steps. The first two network eigenmodes and the projection of this 2DCA –interpreted as a TN– is displayed in Fig. 3. The low-dimensional trajectory is similar to the one found for a stochastic TN with memory (Fig. 2C), as expected since Conway’s game of life is unpredictable, but not totally random.
Figure 4: (A) $10\times 10$ Permutation matrix $\mathbb{P}$. (B) Projection of the TN ${\bf A}(t+1)=\mathbb{P}{\bf A}(t)$ onto the 2d space spanned by the first two network eigenmodes. The trajectory is periodic, mimicking the periodic cycle of $\mathbb{P}$. (C) Spectrum of $\tilde{\mathscr{K}}$ and $\mathbb{P}$. The leading $n$ eigenvalues of $\tilde{\mathscr{K}}$ coincide with those of $\mathbb{P}$, being roots of 1 that yield no growth or decay of dynamic modes, as expected for a permutation operator.
[…]
As a final analysis, we consider a simple sinusoidal TN dynamics where $A_{ij}(t)=\cos(2\pi t/T)+\xi$, with intrinsic period $T=10$ and $\xi\sim N(0,0.1)$ is an extrinsic Gaussian white noise. In this case the dynamics is periodic but the norm of each adjacency matrix severely fluctuates over time. While the POD analysis correctly captures the periodicity in the space spanned by the network eigenmodes, in the approximation to the Koopman operator, a few unstable modes emerge, see the spectrum of $\tilde{\mathscr{K}}$ in Fig. 5A. Now, these (spurious) unstable modes disappear if we conduct an appropriate Takens embedding approach (so as to build a more sophisticated observable $g(\mathscr{X})$), see Fig. 5B where we used a delay dimension $d=10$.
Figure 5: (A) Spectrum of $\tilde{\mathscr{K}}$ associated to a noisy sinusoidal dynamics $A_{ij}(t)=\cos(2\pi t/T)+\xi$. The norm of ${\bf A}(t)$ varies over time, what leads to the emergence of spurious unstable modes in $\tilde{\mathscr{K}}$ (eigenvalues outside the unit circle). (B) Like (A), but after delay-embedding each snapshot $\tilde{\bf a}(t)=[a(t),\dots,a(t-10))]$. This leads to a more sophisticated choice of the observable $g(\mathscr{X})$. In this case the spurious unstable modes disappear.
Discussion – We have proposed two different eigendecompositions applicable to temporal networks, each of them yielding useful information that can be used for a purely data-driven compression (low-dimensional representation) and spectral characterization of the latent graph dynamics, respectively. The former, being a simple extension of POD to the realm of temporal networks, is one out of a suite of possible temporal network embeddings [25, 26] which here exploits linear dimensionality reduction concepts. Further work should explore its applicability across areas including social and biological systems, where (temporal) networks are pervasive. On a technical note, observe that graphs are invariant under node relabelling, whereas adjacency matrices aren’t. This is a limitation, as any a priori obscure node labelling can hinder the patterns observed in the different network eigenmodes, and similarly if node labels are not maintained across network snapshots, the resulting decomposition is less interpretable. Adequate extensions of these methods for unlabelled networks are thus needed [39].
With regard to approximating the Koopman operator, the method provides useful spectral characterization of a linearization of the latent graph dynamics when such dynamics is close to linear and works well when the number of links is approximately constant over time. For TNs where the number of links or the matrix norm severely fluctuates over time, the basic method captures (spurious) unstable dynamic modes that quickly govern the dynamics. In this case, a more accurate dynamical reconstruction is found by delay-embedding the network trajectory. True Koopman mode instabilities emerging in the data-driven linearization of the latent graph dynamics can also benefit from control-theoretic techniques, providing concrete interventions in the temporal network to preserve stability.
Acknowledgments – The author acknowledges partial support from project CSxAI (PID2024-157526NB-I00) and MdM (CEX2021-001164-M) funded by the MICIU/AEI/10.13039/501100011033, and from the European Commission Chips Joint Undertaking project No. 101194363 (NEHIL).